Wednesday November 19
The day shift began with the Antiproton Source (Pbar) stacking, with experts from the Recycler (RR), Tevatron (TeV) and Booster conducting studies, and with MiniBooNE taking beam.
TeV experts put the TeV through a dry squeeze at 10:16 AM and then conducted damper studies until noon.
Preacc experts switched over to the H- Source at 1:23 PM.
Earlier there was a problem that kept Pbar from stacking. At 9:30 PM an expert came in and fixed the problem.
Thursday November 20
The midnight shift began with Pbar stacking, with studies in the TeV and RR, and with MiniBooNE taking beam.
At 1:38 AM, operators began investigating some strange Pbar emittances. They found and replaced a bad power supply for a correction magnet (A:V104). Stacking resumed around 2:30 AM.
Cryo system experts began work on the TeV sector D1 wet engine at 9:30 AM. They changed out a worn belt.
Experts put Pbar into shot setup for transfers of antiprotons into the Main Injector (MI). The shots were successful.
Recycler took beam when available.
Pbar resumed stacking at 5 PM.
At 8 PM, a Pbar expert began a reverse proton study period.
Friday November 21
The midnight shift began with Pbar switching back to stacking mode, with experts conducting TeV and RR studies, and with MiniBooNE taking beam.
By 1 AM, Operations had Pbar ready to resume stacking, but had problems getting beam into the Debuncher. After investigation, operators first found a problem with a kicker (D:IKIK) and then with a Septum (ISEP). Stacking resumed around 2 AM.
Continue with startup. Although the TeV, MI, and Pbar have a lot of work to do yet, the hope is to establish a store this weekend.
Accelerator Update Archive
For Tevatron luminosity charts and the current status of Fermilab's
accelerators and detectors (live!), please go to Fermilab Now
Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: firstname.lastname@example.org.